A battery supplies an EMF of 13 V across a load with a resistance of 22 kΩ. What current passes through the load (in milliAmps)?
Answer #1

Given, V = 13 V. and R = 22 kΩ = 22000 Ω.

Since, I = V/R = (13) / (22000) = 0.00059 A = 0.59 mA (milliAmps)